When a steel block is released from rest on a steep inclined plane, it moves 50.0 cm in 0.404 s. The inclined plane makes an angle of 50.00 with respect to the horizontal. There is a constant force of kinetic friction between the steel block and the inclined surface. What is the kinetic coefficient of the frictional force?
1
Expert's answer
2020-07-21T13:48:51-0400

"mg\sin50°-\mu mg\cos50°=ma"


"S=\frac{at^2}{2}\to a=\frac{2S}{t^2}=\frac{2\cdot0.5}{0.404^2}=6.13m/s^2"


"\mu=\tan50°-\frac{a}{g\cos50°}=\tan50°-\frac{6.13}{9.81\cdot\cos50°}=0.22" Answer